# A Comprehensive Simulation Framework for CXL Disaggregated Memory

### v3

@@ Abstract
Compute eXpress Link (CXL) has emerged as a key enabler of memory disaggregation for future heterogeneous computing systems to expand memory on-demand and improve resource utilization. However, CXL is still in its infancy stage and lacks commodity products on the market, thus necessitating a reliable system-level simulation tool for research and development. In this paper, we propose CXL-DMSim111Open sourced at https://github.com/ferry-hhh/CXL-DMSim., an open-source full-system simulator to simulate CXL disaggregated memory systems with high fidelity at a gem5-comparable simulation speed. CXL-DMSim incorporates a flexible CXL memory expander model along with its associated device driver, and CXL protocol support with CXL.io and CXL.mem. It can operate in both app-managed mode and kernel-managed mode, with the latter using a dedicated NUMA-compatible mechanism. The simulator has been rigorously verified against a real hardware testbed with both FPGA-based and ASIC-based CXL memory prototypes, which demonstrates the qualification of CXL-DMSim in simulating the characteristics of various CXL memory devices at an average simulation error of 4.1%. The experimental results using LMbench and STREAM benchmarks suggest that the CXL-FPGA memory exhibits a 2.88 higher latency than local DDR while the CXL-ASIC latency is 2.18; CXL-FPGA achieves 45-69% of local DDR memory bandwidth, whereas the number for CXL-ASIC is 82-83%. We observed that the performance of CXL memory is 3 more sensitive to Rd/Wr patterns than local DDR, with the max. bandwidth at 74%:26% rather than 50%:50% due to the current compromised CXL+DDR controller design. The study also reveals that CXL memory can significantly enhance the performance of memory-intensive applications, improved by 23 at most with limited local memory for Viper and approximately 16% in bandwidth-sensitive scenarios such as MERCI. Moreover, the simulator’s observability and expandability are showcased with detailed case-studies, highlighting its great potential for research on future CXL-interconnected hybrid memory pool.
I Introduction
@@
With the prevalence of massive data-driven applications such as AI/ML and big data analytics, the demand for larger memory is ever-increasing in today’s heterogeneous parallel computing systems. Over the past two decades, the CPU performance has been boosted dramatically thanks to Moore’s law and multi/many-core scaling. However, the memory capacity and bandwidth per core have been decreasing, which apparently poses a bottleneck for system performance [1]. In modern datacenters, the deployment unit is a monolithic server which contains closely-coupled computing and memory resources. This monolithic architecture for many years is always CPU-biased, leading to memory over-provision across the whole system. It has been observed that more than 50% of the aggregated memory is unused most of the time in production clusters at Google and Facebook [2]. Considering the rising DRAM chip prices in recent years, the under-utilization of memory resources becomes prohibitively expensive, which greatly boosts the TCO of datacenters [3].
@@
With the advent of memory disaggregation technologies, new solutions can be explored to tackle the memory wall and memory under-utilization challenges [4]. Memory disaggregation technologies decouple memory resources from CPUs, providing a feasible option for memory pooling. Conventionally, Remote Direct Memory Access (RDMA) technology is exploited to realize memory disaggregation [2, 5, 6, 7, 8]. But RDMA is based on networking IO semantics which requires specialized NICs and software intervention, leading to a latency multiple orders of magnitude longer than that of local memory access. In recent years, several low-latency and high-bandwidth memory-coherent interconnect protocols arise in industry, which have shown advantages over RDMA for memory disaggregation [9, 10, 11, 12, 13, 14, 15]. Among them, the CXL protocol [16] is very promising and embraced by an increasingly number of semiconductor vendors worldwide. With CXL, memory expansion becomes more flexible over the interconnect fabric while enabling coherent memory access via load/store instructions. Furthermore, the CXL protocol is independent on the underlying memory technology, which can be DRAM, Flash, or even emerging non-volatile memories such as MRAM and RRAM. This facilitates the construction of a unified heterogeneous memory pool for future energy- and cost-efficient computing systems.
@@
Despite its attractive features disclosed in the protocol specifications, CXL is still in its infancy stage and lacks commodity products on the market. As a result, the prior research work on CXL-based memory disaggregation is conducted based on four main methods: software-based emulation [17], software-based simulation, [18, 19, 20], hardware-based emulation [21, 22, 23], and hardware prototyping [9, 24, 25]. However, the software-based emulation such as QEMU fails to model the physical characteristics and internal micro-architecture of real CXL memory devices. The simulation endeavors of gem5-CXL and CXLMemSim are both nascent, with gem5-CXL failing to accurately model the CXL protocol behavior and provide clear access interfaces, while CXLMemSim lacks full-system simulation capabilities and cycle-accurate fidelity, resulting in limited functionality and poor usability. The hardware-based emulation such as remote NUMA lacks CXL protocol support and there is a big difference in the memory access path and performance. As for the fourth method, there are currently no market-ready prototypes of CXL-based memory-disaggregated systems; CXL commodity products are also expensive to produce and purchase. Given the above limitations, there is a clear need for an accurate, cost-effective, and flexible tool for research on CXL-based disaggregated memory systems.
@@
In this paper, we present CXL-DMSim, a full-system CXL Disaggregated Memory Simulator based on gem5 for cycle-accurate simulation, architectural exploration, and evaluation of CXL-interconnected memory systems. CXL-DMSim is as easily configurable as the original gem5 simulator and fits to a variety of CXL devices. It has been rigorously verified and calibrated by a real-world CXL1.1 testbed with both an in-house ASIC CXL memory expander and an FPGA-based CXL device prototype. To the best of our knowledge, CXL-DMSim is the first usable full-system disaggregated memory simulator. The main contributions of this paper are listed below.
@@
A flexible device model of CXL memory expander (Type 3 device) which currently supports both DRAM and Flash as underlying storage media.
@@
Supports for CXL.io and CXL.mem sub-protocols, which are used to enumerate, configure, and access our CXL memory device on CXL-DMSim.
@@
An extensive evaluation that validates CXL-DMSim including performance tests, usability&fidelity tests, real-world app. tests, and observation&expandability tests; this verifies the system’s feasibility and offers guidance for appropriate usage of CXL disaggregated memory.
@@ II-A CXL Protocol
Compute eXpress Link (CXL) is a cache-coherent high-speed interconnect for CPUs, memory expanders, and accelerators [16]. CXL features memory coherence between host memories and device-attached memories, enabling resource pooling and sharing for higher performance and reduced software complexity at lower cost. It was first introduced by Intel in 2019 and since then has been widely embraced by the industry. As of today, CXL protocol has undergone updates with three major versions: 1.0/1.1, 2.0, and 3.0/3.1.
CXL is built on the foundation of the PCIe physical layer. It consists of three sub-protocols, namely CXL.io, CXL.cache, and CXL.mem. CXL.io is functionally equivalent to PCIe in discovering and enumerating devices. CXL.cache enables IO devices and accelerators to directly access and cache the host memory. CXL.mem allows the host to access memories that are attached to CXL devices via load/store instructions.
The three sub-protocols identify three device types for distinct application scenarios. Type 1 devices employ CXL.io and CXL.cache sub-protocols. This type of devices are typically referred to as IO devices such as smart NIC which has no memory attached to it. Type 2 devices (accelerators), such as GPU owning device memory and specialized compute units, utilize all three sub-protocols. Type 3 devices use the CXL.io and CXL.mem sub-protocols and are usually depicted as memory expanders, as illustrated in Fig. 1. This paper focuses on the CXL Type 3 devices for flexible memory expansion and pooling of future heterogeneous computing systems.
@@ II-B1 Software-Based Emulation
QEMU is a widely used and open-source emulator and virtualizer. As an emulator, it can perform full-system emulation which emulates all system components, including the processor and peripherals. In terms of CXL support, QEMU has begun to support CXL 2.0, with a few CXL components such as CXL Host Bridge, CXL Root Ports, CXL Switch and CXL Memory Devices [17]. A recent work named Mess [26] presented a novel memory benchmark which measures bandwidth-latency curves. Based on the curves, a one-size-fit-all memory emulator was proposed to emulate the real-time latency of different types of main memories covering DDR4, DDR5, HBM, and CXL memories.
@@ II-B2 Software-Based Simulation
gem5-CXL [18] is an open-source project we found on GitHub. It crudely adds a custom bus between the original gem5 MemBus and DDR controllers to introduce an extra delay. This extra delay actually models the deteriorated performance of CXL memory in comparison to DDR memory. Another project CXLMemSim [19, 20] is a lightweight trace-driven simulator for the CXL.mem sub-protocol. It allows users to define the desired topology of a modeled memory pool as well as the latency of each component. Basically, it relies on the execution of an unmodified program on a real server. The execution time is divided into multiple epochs, each of which is recalculated in terms of execution time based on the collected memory access events and the modeled latency of CXL memory.
@@ II-B3 Hardware-Based Emulation
Due to the unavailability of commercial CXL hardware, many researchers have conducted emulations of CXL-expanded memory exploiting the NUMA mechanism. These studies [21, 22, 23] leverage NUMA-enabled hardware to emulate CXL disaggregated memory without directly handling the CXL protocol. Given the similar characteristics in bandwidth and latency between CXL memory and remote NUMA memory, this approach provides a reference for the study of computing systems with real CXL memory.
@@ II-C Motivation
Although QEMU serves as a generic system emulator, it is primarily used for functional-level emulation and verification. It does not model the physical characteristics and internal micro-architecture of CXL devices. Thus, QEMU is unsuitable for architectural exploration and accurate performance evaluation of CXL-based systems. Similarly, the Mess emulator targets at reproducing the bandwidth-latency curves of the SystemC model of a Micron’s CXL memory expander; Real CXL memory devices may behave differently in a CXL-connected system. We also argue that the cpuBW value calculated from the most recently 1000 memory operations () may not be able to accurately capture the realistic memory bandwidth (messBW) under real-world applications.
Additionally, existing software-based simulation attempts are far away from readiness. gem5-CXL connects a simulated CXL device to the MemBus rather than mounting it on the IOBus to simulate peripherals. It does not model the behaviors of the CXL protocol (based on PCIe PHY) and lacks clear access interfaces with the modeled device. The project has not been updated for four years and is currently in a stagnant state. CXLMemSim is not a full-system simulator in essence, as neither the CXL protocol and system architecture nor the associated software stack are modeled. It lacks cycle accuracy and has limited functionalities, resulting in a poor usability.
As for the other emulation approach, the remote NUMA memory lacks realism and accuracy in mimicking CXL memory. Specifically, there is a significant difference in access latency and bandwidth between these two types of memory, as emphasized in [25]; they also have different memory access paths (see Fig. 1). In terms of the hardware prototype method, there are currently no market-ready hardware prototypes of CXL-based memory disaggregation systems. Existing hardware prototypes and commercial devices are expensive to produce, due to the long silicon development cycle. Moreover, hardware prototypes also lack flexibility in system configuration and observation.
@@ III-A Simulator Architecture
Fig. 2 shows the overall architecture of CXL-DMSim, with three main components built on top of gem5. First, we added a CXL memory expander model, illustrated as CXL Host-managed Device Memory (HDM) in the figure, to the gem5 simulator. It connects to the system as a PCI device with standalone memory space through IOBus (see Fig. 3). Second, we implemented the CXL sub-protocols CXL.io and CXL.mem, which are required for Type 3 devices. The CXL.io sub-protocol is used for the CPU to enumerate and configure CXL devices; it is achieved by reusing the original PCI protocol in gem5. The CXL.mem sub-protocol allows the CPU to access the CXL HDM directly; it is achieved by integrating new CXL.mem packets as well as by extending internal components such as bridge in gem5.
Third, we designed a dedicated device driver and a NUMA-compatible memory management mechanism in the guest OS, to manage the allocation and deallocation of CXL HDM as well as to provide interfaces to upper-level user applications.
Depending on the CXL HDM management mechanism in the guest OS, CXL-DMSim provides users with two ways to use our CXL memory expander: Application Managed (AM) and Kernel Managed (KM). In the AM mode, applications can allocate and free CXL HDM using the memkind library or the mmap system call interface provided by our device driver. This usage mode offers users more flexibility and granularity in managing CXL HDM but also imposes the pressure of modifying legacy programs. In the KM mode, the CXL memory expander is exposed to the OS kernel as a CPU-less (or memory-only) NUMA node, with the kernel managing the allocation and release of CXL HDM transparently to applications. Users can seamlessly utilize CXL HDM through existing tools such as numactl. However, it is evident that in this mode the difference from the DDR memory is only reflected in the “NUMA distance”. Users are not able to perceive and exploit the unique characteristics (e.g., persistence, large volume, and low power consumption) of different memory technologies behind the CXL interface. Note that the KM mode is consistent with the mainstream usage of CXL memory expanders on real-world hardware platforms.

### v6

# CXL-DMSim: A Full-System CXL Disaggregated Memory Simulator With Comprehensive Silicon Validation

@@ Abstract
Compute eXpress Link (CXL) has emerged as a key enabler of memory disaggregation for future heterogeneous computing systems to expand memory on-demand and improve resource utilization. However, CXL is still in its infancy stage and lacks commodity products on the market, thus necessitating a reliable system-level simulation tool for research and development. In this paper, we propose CXL-DMSim111Open sourced at https://github.com/ferry-hhh/CXL-DMSim., an open-source full-system simulator to simulate CXL disaggregated memory systems with high fidelity at a gem5-comparable simulation speed. CXL-DMSim incorporates a flexible CXL memory expander model along with its associated device driver, and CXL protocol support with CXL.io and CXL.mem. It can operate in both app-managed mode and kernel-managed mode, with the latter using a dedicated NUMA-compatible mechanism. The simulator has been rigorously verified against a real hardware testbed with both FPGA- and ASIC-based CXL memory devices, which demonstrates the qualification of CXL-DMSim in simulating the characteristics of various CXL memory devices at an average simulation error of 3.4%. The experimental results using LMbench and STREAM benchmarks suggest that the CXL-FPGA memory exhibits a 2.88 higher latency than local DDR while the CXL-ASIC latency is 2.18; CXL-FPGA achieves 45-69% of local DDR memory bandwidth, whereas the number for CXL-ASIC is 82-83%. The study also reveals that CXL memory can significantly enhance the performance of memory-intensive applications, improved by 23 at most with limited local memory for Viper key–value database and approximately 60% in memory-bandwidth-sensitive scenarios such as MERCI. Moreover, the simulator’s observability and expandability are showcased with detailed case-studies, highlighting its great potential for research on future CXL-interconnected hybrid memory pool.
Index Terms:
Compute express link, Memory disaggregation, Full-system simulator, Hardware prototyping, Performance benchmarking.I Introduction
@@
With the prevalence of massive data-driven applications such as AI/ML and big data analytics, the demand for larger memory is ever-increasing in today’s heterogeneous parallel computing systems. Over the past two decades, the CPU performance has been boosted dramatically thanks to Moore’s law and multi/many-core scaling. However, the memory capacity and bandwidth per core have been decreasing, which apparently poses a bottleneck for system performance [1]. In modern datacenters, the deployment unit is typically a monolithic server which contains closely-coupled computing and memory resources. This monolithic architecture for many years is always CPU-biased, leading to memory over-provision across the whole system. It has been observed that approximately 40% of aggregated memory is unused most of the time in the hyperscale infrastructure at Azure and Google, primarily due to stranded memory (leftover memory when all cores of a server are rented out) and untouched memory (rented yet not actually used) [2, 3]. Considering the rising DRAM chip price in recent years, the under-utilization of memory resources becomes prohibitively expensive, which greatly boosts the TCO of today’s datacenters [2].
@@
With the advent of memory disaggregation technologies, new solutions can be explored to tackle the memory wall and memory under-utilization challenges [4]. Memory disaggregation technologies decouple memory resources from CPUs, providing a feasible option for memory pooling. Conventionally, remote direct memory access (RDMA) technology is exploited to realize memory disaggregation [5, 6, 7]. But RDMA is based on networking IO semantics which requires specialized NICs and software intervention, leading to a latency multiple orders of magnitude longer than that of local memory access. In recent years, several low-latency and high-bandwidth cache-coherent interconnect protocols such as OpenCAPI, CCIX, CXL arise in industry, which have shown advantages over RDMA for memory disaggregation [8, 9, 10]. Among them, the CXL protocol [11] is considered most promising and embraced by an increasingly number of semiconductor vendors worldwide. With CXL, memory expansion becomes more flexible over the interconnect fabric while enabling coherent memory access via load/store instructions. Furthermore, the CXL protocol is independent on the underlying memory technology, which can be DRAM, Flash, or even emerging non-volatile memories such as MRAM and RRAM. This facilitates the construction of a unified heterogeneous memory pool for future energy- and cost-efficient computing systems.
@@
Despite its attractive features disclosed in the protocol specifications, CXL is still in its infancy stage and lacks commodity products on the market. As a result, the prior research work on CXL-based memory disaggregation is conducted based on four main methods: software-based emulation [12, 13], software-based simulation [14, 15, 16, 17], hardware-based emulation [18, 19], and hardware prototyping [20, 21, 22]. However, the software-based emulation such as QEMU fails to model the physical characteristics and internal micro-architecture of real CXL memory devices. The simulation attempts of gem5-CXL and CXLMemSim are both immature and flawed; gem5-CXL fails to accurately model the CXL protocol behavior and to provide clear access interfaces, while CXLMemSim lacks full-system simulation capabilities and cycle-level fidelity resulting in limited functionality and poor usability. The hardware-based emulation such as remote NUMA lacks CXL protocol support and there is a big difference in the memory access path and performance. As for the fourth method, there are currently no market-ready prototypes of CXL-based memory-disaggregated systems; CXL commodity products are also expensive to produce and purchase. Given the above limitations, there is a clear need for an accurate, cost-effective, and flexible tool for research on CXL-based disaggregated memory systems.
@@
In this paper, we present CXL-DMSim, a full-system CXL Disaggregated Memory Simulator for cycle-level simulation, architectural exploration, and evaluation of CXL-interconnected memory systems. CXL-DMSim is as easily configurable as the original gem5 simulator and fits to a variety of CXL devices. It has been rigorously verified and calibrated against a real-world CXL 1.1 testbed with both an in-house ASIC-based CXL memory expander and an FPGA-based CXL device prototype. To the best of our knowledge, CXL-DMSim is the first usable full-system disaggregated memory simulator. The main contributions of this paper are listed below.
@@
A flexible device model of CXL memory expander (type-3 device) which currently supports DRAM, NVM and Flash as underlying storage media.
@@
Currently CXL1.1+ supports for CXL.io and CXL.mem sub-protocols, which are used to enumerate, configure, and access our CXL memory device on CXL-DMSim.
@@
An extensive silicon measurement and evaluation that validates CXL-DMSim including performance tests, usability&fidelity tests, real-world application tests, and observation&expandability tests; this verifies the system’s feasibility and offers guidelines for appropriate usage of CXL disaggregated memory.
@@ II-A CXL Protocol
Compute eXpress Link (CXL) is a cache-coherent high-speed interconnect for CPUs, memory expanders, and accelerators [11]. CXL features memory coherence between host memories and device-attached memories, enabling resource pooling and sharing for higher performance and reduced software complexity at lower cost. It was first introduced by Intel in 2019 and since then has been widely embraced by the industry. As of today, CXL has undergone updates with three major versions: 1.0/1.1, 2.0, and 3.0/3.1/3.2.
CXL is built on the PCIe physical layer and includes three sub-protocols: CXL.io, CXL.cache, and CXL.mem. CXL.io is functionally equivalent to PCIe in discovering and enumerating devices. CXL.cache enables I/O devices and accelerators to access and cache the host memory. CXL.mem allows the host to access memories that are attached to CXL devices via load/store instructions. The three sub-protocols identify three device types for distinct application scenarios. Type-1 devices employ CXL.io and CXL.cache sub-protocols. This type of devices are typically referred to as I/O devices such as SmartNIC which has no memory attached to it. Type-2 devices (accelerators), such as GPU with dedicated device memory and specialized compute units, utilize all three sub-protocols. Type-3 devices use the CXL.io and CXL.mem sub-protocols and are usually depicted as memory expanders, including both direct-attached and switch-attached ones, as shown in Fig. 1. This paper focuses on type-3 devices for flexible memory expansion and pooling of future heterogeneous computing systems.
@@ II-B1 Software-Based Emulation
QEMU is a widely used and open-source emulator and virtualizer. As an emulator, it can perform full-system emulation which emulates all system components, including the processor and peripherals. In terms of CXL support, QEMU has begun to support CXL 2.0, with a few CXL components such as the CXL host bridge, CXL root ports, CXL switch, and CXL memory devices [12]. A recent work named Mess [13] presented a novel memory benchmark which measures bandwidth-latency curves. Based on the curves, a one-size-fit-all memory emulator was proposed to emulate the real-time latency of different types of main memories covering DDR4, DDR5, HBM, and CXL memories.
@@ II-B2 Software-Based Simulation
gem5-CXL [14] is an open-source project which crudely adds a custom bus between the original gem5 MemBus and DDR controllers to introduce an extra delay. This extra delay actually models the deteriorated performance of CXL memory in comparison to DDR memory. DisaggSim [15] is similar to gem5-CXL but it does a step further by modeling a large-scale disaggregated memory system with Garnet model forming an interconnect fabric. DRackSim [16] is an application-level simulation framework, which uses PinTool at the front-end to collect instruction traces from applications running on real machines. Its back-end performs cycle-level simulation based on these traces. CXLMemSim [17] is also a trace-driven simulator. Users can specify memory pool topologies and component latencies. It runs unmodified applications on real servers, divides execution into epochs, and recalculates execution time per epoch based on recorded memory accesses and modeled CXL latency.
@@ II-B3 Hardware-Based Emulation
Due to the unavailability of commercial CXL hardware, many researchers have conducted emulations of CXL-expanded memory exploiting the NUMA mechanism. These studies [18, 19] leverage NUMA-enabled hardware to emulate CXL memory without directly handling the CXL protocol. Given the similar characteristics in bandwidth and latency between CXL memory and remote NUMA memory, this approach provides a reference for the study of computing systems with real CXL memory.
@@ II-C Motivation
Although QEMU serves as a generic system emulator, it is primarily used for functional-level emulation and verification. It does not model the physical characteristics and internal micro-architecture of CXL devices. Thus, QEMU is unsuitable for architectural exploration and accurate performance evaluation of CXL-based systems. Similarly, the Mess emulator targets at reproducing the bandwidth-latency curves of the SystemC model of a Micron’s CXL memory expander; real CXL memory devices may behave differently in a CXL-connected system. Moreover, Mess also lacks modeling of the CXL protocol, which limits its capability for architectural exploration especially when considering advanced CXL features such as memory pooling and sharing.
Additionally, existing software-based simulation attempts are far away from readiness. gem5-CXL connects a simulated CXL device to the MemBus rather than mounting it on the IOBus to simulate peripherals. It does not model the behaviors of the CXL protocol (based on PCIe PHY) and lacks clear access interfaces with the modeled device. The project has not been updated for four years and is currently in a stagnant state. DisaggSim and DRackSim perform coarse-grained modeling of CXL-enabled systems by abstracting key behaviors and modeling generalized components of disaggregated memory. However, these abstractions and simplifications introduce significant simulation errors. Specifically, DisaggSim lacks calibration, whereas DRackSim is calibrated only against gem5 and public DirectCXL data, showing up to 12% average error. Worse still, DRackSim’s custom host model runs 2 slower than gem5. CXLMemSim is not a full-system simulator in essence, as neither the CXL protocol and system architecture nor the associated software stack are modeled. It lacks cycle accuracy and has limited functionalities, resulting in poor usability.
As for the hardware-based emulation approach, the remote NUMA memory lacks realism and accuracy in mimicking CXL memory. Specifically, there is a significant difference in access latency and bandwidth between these two types of memory, as emphasized in [22]; they also have different memory access paths (see Fig. 1). In terms of the hardware prototype method, there are currently no market-ready hardware prototypes of CXL-based memory disaggregation systems. Existing hardware prototypes and commercial devices are expensive to produce, due to the long silicon development cycle. Moreover, hardware prototypes also lack flexibility in system configuration and observation.
@@ III-A Simulator Architecture
Fig. 2 shows the overall architecture of CXL-DMSim, with three main components built on top of gem5. First, we added a CXL memory expander model, illustrated as CXL host-managed device memory (HDM) in the figure, to the gem5 simulator. Second, we implemented the CXL sub-protocols CXL.io and CXL.mem, which are required for type-3 devices. The CXL.io sub-protocol is used for the CPU to enumerate and configure CXL devices; it is achieved by reusing the original PCI protocol in gem5. The CXL.mem sub-protocol allows the CPU to access the CXL HDM directly; it is achieved by integrating new CXL.mem packets as well as by extending internal components such as bridge in gem5. Third, we designed a dedicated device driver and a NUMA-compatible memory management mechanism in the guest OS, to manage the allocation and deallocation of CXL HDM as well as to provide interfaces to upper-level user applications.
Depending on the CXL HDM management mechanism in the guest OS, CXL-DMSim provides users with two ways to use our CXL memory expander: application-managed (AM) and kernel-managed (KM). In the AM mode, applications can allocate and free CXL HDM using the memkind library or the mmap system call interface provided by our device driver. This usage mode offers users more flexibility and granularity in managing CXL HDM but also imposes the pressure of modifying legacy programs. In the KM mode, the CXL memory expander is exposed to the OS kernel as a CPU-less NUMA node, with the kernel managing the allocation and release of CXL HDM transparently to applications. Users can seamlessly utilize CXL HDM through existing tools such as numactl. However, it is evident that in this mode the difference from the DDR memory is only reflected in the “NUMA distance”. Users are not able to perceive and exploit the unique characteristics (e.g., persistence, large volume, and low power consumption) of different memory technologies behind the CXL interface. Note that the KM mode is consistent with the mainstream usage of CXL memory expanders on real-world hardware platforms.